OpenStudy (anonymous):

Quadratic equations: Solve for x: X4 - 5X2 + 4 = 0

OpenStudy (anonymous):

of course

OpenStudy (campbell_st):

look at the equation this way \[(x^2)^2 - 5(x^2) + 4 = 0\] or there is a method of substitution where you let \[u = x^2\] so you need to solve \[u^2 - 5u + 4 = 0\] both equations can be solved by factoring If you use the equation involving u, when you get the solutions substitute x^2 back in and then solve for x hope this helps

OpenStudy (anonymous):

why do I have to use x^2?

OpenStudy (campbell_st):

well the equation you have been given falls into the group called equations that are reducible to quadratics... so instead of having \[ax^2 + bx + c = 0\] you have \[a(x^2)^2 + b(x^2) + c = 0\] or it can be written as \[au^2 + bu + c = 0\] when you get it into a quadratic form it can be solved using factoring, completing the square or general quadratic formula

OpenStudy (anonymous):

um I still don't see why x^2 was used couldn't we have used X^4?

OpenStudy (campbell_st):

well that of it this way using index laws \[(x^2)^2 = x^{2 \times 2} = x^4\]

OpenStudy (campbell_st):

and no you can use the generarl quadratic formula on an equation that is a quartic... you need to get it to a quadratic form.... and there will be 4 solutions... and the equation is quite simple to factor...

OpenStudy (anonymous):

well in my book that is the example of how to solve one of these things and I guess I have t know this way for the test.

OpenStudy (anonymous):

can you finish the problem?

OpenStudy (campbell_st):

this is the factorization you can have \[x^4 - 5x^2 + 4 = (x^2 -4)(x^2 -1) \] or by making a substitution \[u = x^2\] the factored from is \[u^2 - 5u + 4 = (u -4)(u - 1)\] so in either case after factoring you need to solve \[x^2 - 4 = 0..... and......x^2 - 1 = 0\]

OpenStudy (campbell_st):

thats how you do it.... and you'll find there are 4 solutions.

OpenStudy (anonymous):

alright thanks!